OpenStudy (help!!!!):

2. If Sally can paint a house in 4 hours, and John can paint the same house in 6 hour, how long will it take for both of them to paint the house together?

OpenStudy (anonymous):

2 just subtract the times from each other

OpenStudy (help!!!!):

what would be the exact number

OpenStudy (anonymous):

2

OpenStudy (help!!!!):

the answer is 2 hours and 24 minutes. i dont know how to get the answer

OpenStudy (anonymous):

If Sally can paint the house in 4 hours, then in 1 hour she can paint 1/4 of the house. If John can paint the house in 6 hours, then in 1 hour he can paint 1/6 of the house. Let x = number of hours it would take them together to paint the house, then working together, in 1 hour they can paint 1/x of the house.

OpenStudy (help!!!!):

can you show me the step by chance?

OpenStudy (anonymous):

simple here's the formula for it

OpenStudy (help!!!!):

thath would be 10/24. What would be next?

OpenStudy (anonymous):

thanks for the medal :)

OpenStudy (help!!!!):

ur welcome and thx 4 the help

OpenStudy (anonymous):

no problem

OpenStudy (help!!!!):

can you be one of my fans?

OpenStudy (help!!!!):

can you be one of my fans?

OpenStudy (anonymous):

I am now a fan:)

OpenStudy (help!!!!):

ty bye

OpenStudy (raden):

multiply of them divided by addition (4*6)/(4+6) = 24/10 = 2.4 h = 2h + 0.4 h = 2h + 0.4*60 m = 2h + 24m